OpenStudy (anonymous):

*Composition of Functions* f(x)=x^2 g(x) = sqrt(x-2) what is g o f(x) and g o g(x)?????

OpenStudy (anonymous):

gof(x)= sqrt( x^2-2)

OpenStudy (anonymous):

gog(x) = sqrt{( sqrt(x-2)-2}

OpenStudy (anonymous):

I know g o f(x) = \[\sqrt{x^2-2}\] and then the -2 you can just rewrite as \[i \sqrt{2?}\] and pull out the x or no?

OpenStudy (anonymous):

you can't pull out x

OpenStudy (anonymous):

ah okay I thought for some reason you could since it is being squared within a square root sign

OpenStudy (anonymous):

gog(x)= \[\sqrt{\sqrt{x-2}-2}\]

OpenStudy (anonymous):

that is not possible

OpenStudy (anonymous):

0.0 yikes....how about the first one, what would my domain be? wolfram alpha saying it would be x >= -2

OpenStudy (anonymous):

sqrt(-2)**

OpenStudy (anonymous):

in first case domain must b greater or equal to 2